a caterer os planning a large dinner party for 1001 guests. Each table must seat and equal number of guests, more than 1 guest and fewer than 20. Determine the number of possible table sizes. If the caterers use 11 person tables, how many tables will they need?
Answer by kevwill(135)

The prime factorization of 1001 is 7*11*13, so with the constraint that the tables seat between 2 and 19 guests, we can use tables that seat either 7, 11, or 13 guests.
For tables that seat 7, the caterer would need 11*13 = 143 tables.
For tables that seat 11, the caterer would need 7*13 = 91 tables.
For tables that seat 13, the caterer would need 7*11 = 77 tables.